OpenStudy (anonymous):

2(tan)^2 45 degrees - 3 tan 60 deg. + cot 30 deg. + 1 will fan and medal

OpenStudy (anonymous):

so if substituted, 2 (1)^2 - 3 (sqr. of 3) + sqr. of 3 + 1

OpenStudy (thomas5267):

\[2\tan^2(45^\circ)-3\tan(60^\circ)+\cot(30^\circ)+1?\]

OpenStudy (anonymous):

yep

OpenStudy (eyust707):

Simplify?

OpenStudy (anonymous):

yeah

OpenStudy (thomas5267):

\[ 2(1)^2-3\sqrt{3}+\sqrt{3}+1=3-2\sqrt{3} \] I think the answer in this form should be okay